Dwumiany

1. Wykaż, że dla k, n ∈ N, takich, że 0 k n zachodzi: \binom{n}{k} = \binom{n}{n−k}

2. Wykaż, że dla k, n ∈ N, takich, że 0 k n zachodzi: \binom{n}{k} = \binom{n−1}{k} + \binom{n−1}{k−1}.

3. Wykaż, że dla k, n ∈ N, takich, że 0 k n zachodzi: \binom{n+1}{k+1} = \binom{n}{k} \frac{n+1}{k+1} .

4. Wykaż, że dla n ∈ N zachodzi: ∑ _{i=0}^{n} \binom{n}{i} = 2^{n}.

5. Wykaż, że dla n ∈ N zachodzi: ∑ _{i=0}^{n} (−1)^{i} \binom{n}{i} = 0.

6. Wykaż, że dla n ∈ N zachodzi: ∑ _{i=0}^{n} 2^{i} \binom{n}{i} = 3^{n}.

7. Wykaż, że dla n ∈ N zachodzi: ∑ _{i=0}^{n} i\binom{n}{i} = n2^{n−1} .

8. Wykaż, że dla n ∈ N zachodzi: ∑ _{i=0}^{n} i(i − 1)\binom{n}{i} = n(n − 1)2^{n−2} .

9. Wykaż, że dla n ∈ N zachodzi: ∑ _{i=0}^{n} \binom{n}{i} ^{2} = \binom{2n}{n}.

10. Wykaż, że dla j, k, n ∈ N takich, że n + k j zachodzi: ∑ _{i=0}^{n} \binom{n}{i} \binom{k}{j−i} = \binom{n+k}{j}.

11. Wykaż, że dla n, m ∈ N takich, że m n zachodzi: ∑ _{i=n}^{m} \binom{i}{n} = \binom{m+1}{n+1}.